# Iterative Data-Consistent Inversion with Multiple Push-forward Constraints

### v1

@@ 6.2.2 Iterative Results
Figures 12 and 13 summarize the QoI marginals associated with the first and last iterations, respectively, of the first epoch. In the plots of these figures, the blue dotted curves are the marginals associated with the push-forward of the updated distributions while the orange curves are the marginals associated with the observed distribution. In Figure 12, the two marginals associated with $Q_{1}$ (top left plot) lay almost perfectly on top of each other while significant visual discrepancies are evident between the two marginal densities associated with all of the remaining QoI. This is expected since the only push-forward constraint utilized at this point is the one associated with $Q_{1}$. However, once all of the push-forward constraints are utilized by iteration #12, the two marginals for all of the QoI appear qualitatively similar as seen in Figure 13. Despite the qualitative similarities, there are some clear numerical discrepancies between several of these marginal densities. For instance, the marginals associated with $Q_{5}$ (shown in the middle plot of the second row) and with $Q_{9}$ (shown in the right plot of the third row) have the most visually evident discrepancies. Further epochs are clearly required before numerical convergence, as measured by the absolute KL tolerance of ${1}\mathrm{E}{-6}$, is reached.
@@
The algorithm terminated at the third epoch due to the absolute KL divergence stopping criteria. In other words, the KL divergence of every QoI marginal associated with the push-forward of the updated distribution at the end of this epoch was within ${1}\mathrm{E}{-6}$ of the associated observed marginal. This is visually confirmed by the array of QoI marginal plots shown in Figure 15 that is arranged in the same order as Figures 12 and 13.
@@ 7 Conclusions and Future Work
  • Optimization of Initial Distribution: The final converged solution ($P^{\infty}$) is dependent upon the choice of the initial measure ($P^{0}=P_{\text{init}}$). Future work will explore strategies for optimizing the choice of $P^{0}$ to improve inference speed or target specific features of the data-generating distribution. This includes analyzing the convergence properties when using a sequence of varying initial distributions that are mutually absolutely continuous after the first epoch. This also includes analyzing what conditions on $P_{\text{init}}$ are necessary to guarantee $P_{\text{init}}\ll P_{\text{up}}$ to conclude optimality of the backward minimization problem.

### v2

@@ 6.2.2 Iterative Results
Figures 15 and 16 summarize the QoI marginals associated with the first and last iterations, respectively, of the first epoch. In the plots of these figures, the blue dotted curves are the marginals associated with the push-forward of the updated distributions while the orange curves are the marginals associated with the observed distribution. In Figure 15, the two marginals associated with $Q_{1}$ (top left plot) lay almost perfectly on top of each other while significant visual discrepancies are evident between the two marginal densities associated with all of the remaining QoI. This is expected since the only push-forward constraint utilized at this point is the one associated with $Q_{1}$. However, once all of the push-forward constraints are utilized by iteration #12, the two marginals for all of the QoI appear qualitatively similar as seen in Figure 16. Despite the qualitative similarities, there are some clear numerical discrepancies between several of these marginal densities. For instance, the marginals associated with $Q_{5}$ (shown in the middle plot of the second row) and with $Q_{9}$ (shown in the right plot of the third row) have the most visually evident discrepancies. Further epochs are clearly required before numerical convergence, as measured by the absolute KL tolerance of ${1}\mathrm{E}{-6}$, is reached.
@@
The algorithm terminated at the third epoch due to the absolute KL divergence stopping criteria. In other words, the KL divergence of every QoI marginal associated with the push-forward of the updated distribution at the end of this epoch was within ${1}\mathrm{E}{-6}$ of the associated observed marginal. This is visually confirmed by the array of QoI marginal plots shown in Figure 18 that is arranged in the same order as Figures 15 and 16.
@@ 7 Conclusions and Future Work
  • Optimization of Initial Distribution: The final converged solution ($P^{\infty}$) is dependent upon the choice of the initial measure ($P^{0}=P_{\text{init}}$). Future work will explore strategies for optimizing the choice of $P^{0}$ to target specific features of the data-generating distribution. This includes analyzing the convergence properties when using a sequence of varying initial distributions that are mutually absolutely continuous after the first epoch. This also includes analyzing what conditions on $P_{\text{init}}$ are necessary to guarantee $P_{\text{init}}\ll P_{\text{up}}$ to conclude optimality of the backward minimization problem.